# Interplay of orbital-selective Mott criticality and flat-band physics in La3Ni2O6

### v3

@@ Abstract
Figure 3: Temperature-dependent Ni-$e_{g}$ data of La3Ni2O6 from DFT+sicDMFT. Left: Low-energy spectral function, and right: respective imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$.
@@
At first glance, the low-energy spectral weight at $T=50$ K seemingly shows a sharp quasiparticle(QP)-like peak, but the detailed resolution (see inset of upper panel of Fig. 2a) displays a narrow (pseudo)gap feature of $\sim$50 meV size, in very good agreement with the experimental semiconducting gap of La3Ni2O6 [40]. From the local spectral function, this (pseudo)gap is exclusively linked to Ni-$d_{z^{2}}$, while Ni-$d_{x^{2}-y^{2}}$ is Mott-gapped orbital selectively on a larger scale of $\sim 0.85$ eV. Concerning the local-orbital fillings, the value $n_{x^{2}-y^{2}}=1.10$ marks the expected near-half-filled scenario for Mott-gapped in-plane $d_{x^{2}-y^{2}}$, whereas the higher value $n_{z^{2}}=1.27$ is realized for out-of-plane $d_{z^{2}}$. As shown in the left panel of Fig. 3, the intriguing low-energy spectrum results from a metal-to-insulator transition with lowering temperature. While the Ni-$d_{x^{2}-y^{2}}$ orbital is Mott-gapped already well above 100 K, the Ni-$d_{z^{2}}$ still displays metallic character with a seemingly QP-like peak at $T=150$ K. But in the spectrum at $T=100$ K, the opening of a (pseudo)gap for the latter orbital starts to become visible. Concomitantly, the imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$, shown in the right panel of Fig. 3, looks Fermi-liquid like for Ni-$d_{z^{2}}$ over a wide frequency range. Yet it bends downwards for smallest $\omega_{n}$ at $T=50$ K, i.e. signaling the established (pseudo)gap feature. The larger Ni-$d_{x^{2}-y^{2}}$ self-energy also changes qualitative behavior in the (pseudo)-gap regime, pointing towards an intricate coupling scenario between both Ni-$e_{g}$ sectors.
@@
Figures 2b,c display the k-resolved spectral function for La3Ni2O6. The original Ni-$t_{2g}$ DFT flat-band is repelled from the Fermi level by correlations. Instead, the Ni-$d_{z^{2}}$ flat-band is shifted to lowest energy. This correlated Ni-$d_{z^{2}}$ flat band becomes rather incoherent and splits, giving rise to the small (pseudo)gap at low energy. Hence, flat-band physics as observed in La3Ni2O7 is also present in La3Ni2O6, yet with a different twist. The substantial Ni-$e_{g}$ orbital anisotropy in the reduced bilayer renders $d_{x^{2}-y^{2}}$ closer to Mottness, whereas $d_{z^{2}}$ is distant from Mott criticality. But due to its low-energy flat-band electrons scattering with Mott-localized $d_{x^{2}-y^{2}}$ electrons, it still becomes (pseudo)gapped. The question about an intuitive explanation for this Ni-$d_{z^{2}}$ (pseudo)gapping arises. It should be based on the competition between residual (flat-band) hopping and local-moment formation. Because of the already localized Ni-$d_{x^{2}-y^{2}}$ electron with $S=1/2$, the system may gain exchange energy from hybridizing with flat-band Ni-$d_{z^{2}}$ electrons. It has in fact been shown [62], that such a two-band orbital-selective regime is described by a ferromagnetic Kondo-lattice Hamiltonian. Non-Fermi-liquid behavior and pseudogap features are possible solutions of such ferromagnetic/underscreened Kondo-lattice problems [62, 63, 64, 65]. Here, the scattering between localized Ni-$d_{x^{2}-y^{2}}$ and flat-band Ni-$d_{z^{2}}$ leads to a strong (pseudo)gap.

### v4

@@ Abstract
Figure 3: Effect of temperature on the Ni-$e_{g}$ states of La3Ni2O6 from DFT+sicDMFT. (a) Temperature dependence of low-energy spectral function (left), and respective imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$ (right). (b) Fatspec representation of the k-resolved spectral function $A({\bf k},\omega)$ for $T=150$ K.
@@
At first glance, the low-energy spectral weight at $T=50$ K seemingly shows a sharp quasiparticle(QP)-like peak, but the detailed resolution (see inset of upper panel of Fig. 2a) displays a narrow (pseudo)gap feature of $\sim$50 meV size, in very good agreement with the experimental semiconducting gap of La3Ni2O6 [40]. From the local spectral function, this (pseudo)gap is exclusively linked to Ni-$d_{z^{2}}$, while Ni-$d_{x^{2}-y^{2}}$ is Mott-gapped orbital selectively on a larger scale of $\sim 0.85$ eV. Concerning the local-orbital fillings, the value $n_{x^{2}-y^{2}}=1.10$ marks the expected near-half-filled scenario for Mott-gapped in-plane $d_{x^{2}-y^{2}}$, whereas the higher value $n_{z^{2}}=1.27$ is realized for out-of-plane $d_{z^{2}}$. As shown in the left panel of Fig. 3a, the intriguing low-energy spectrum results from a metal-to-insulator transition with lowering temperature. While the Ni-$d_{x^{2}-y^{2}}$ orbital is Mott-gapped already well above 100 K, the Ni-$d_{z^{2}}$ still displays metallic character with a seemingly QP-like peak at $T=150$ K. But in the spectrum at $T=100$ K, the opening of a (pseudo)gap for the latter orbital starts to become visible. For $T=150$ K, Fig. 3b indeed exhibits the semi-coherent Ni-$d_{z^{2}}$ dominated flat band very close to the Fermi level. Concomitantly, the imaginary part of the Matsubara self-energy $\Sigma(i\omega_{n})$, shown in the right panel of Fig. 3a, looks Fermi-liquid like for Ni-$d_{z^{2}}$ over a wide frequency range. Yet it bends downwards for smallest $\omega_{n}$ at $T=50$ K, i.e. signaling the established (pseudo)gap feature. The larger Ni-$d_{x^{2}-y^{2}}$ self-energy also changes qualitative behavior in the (pseudo)-gap regime, pointing towards an intricate coupling scenario between both Ni-$e_{g}$ sectors.
@@
Figures 2b,c display the k-resolved spectral function for La3Ni2O6 at T=50KT=50\,K. The original Ni-$t_{2g}$ DFT flat-band is repelled from the Fermi level by correlations. Instead, the Ni-$d_{z^{2}}$ flat-band is shifted to lowest energy. This correlated Ni-$d_{z^{2}}$ flat band becomes rather incoherent and splits, giving rise to the small (pseudo)gap at low energy. Hence, flat-band physics as observed in La3Ni2O7 is also present in La3Ni2O6, yet with a different twist. The substantial Ni-$e_{g}$ orbital anisotropy in the reduced bilayer renders $d_{x^{2}-y^{2}}$ closer to Mottness, whereas $d_{z^{2}}$ is distant from Mott criticality. But due to its low-energy flat-band electrons scattering with Mott-localized $d_{x^{2}-y^{2}}$ electrons, it still becomes (pseudo)gapped. The question about an intuitive explanation for this Ni-$d_{z^{2}}$ (pseudo)gapping arises. It should be based on the competition between residual (flat-band) hopping and local-moment formation. Because of the already localized Ni-$d_{x^{2}-y^{2}}$ electron with $S=1/2$, the system may gain exchange energy from hybridizing with flat-band Ni-$d_{z^{2}}$ electrons. It has in fact been shown [62], that such a two-band orbital-selective regime is described by a ferromagnetic Kondo-lattice Hamiltonian. Non-Fermi-liquid behavior and pseudogap features are possible solutions of such ferromagnetic/underscreened Kondo-lattice problems [62, 63, 64, 65]. Here, the scattering between localized Ni-$d_{x^{2}-y^{2}}$ and flat-band Ni-$d_{z^{2}}$ leads to a strong (pseudo)gap.